can you explain how to do this problem x+26=2x+1(x+4)
Answer by stanbon(75887)

x+26=2x+1(x+4)
---
x + 26 = 2x + x+4
---
x + 26 = 3x+4
---
22 = 2x
x = 11
============
Cheers,
Stan H.